Alex earns $55,000 as a salesman. he also earns $100 per a month if he reaches his sales goal. Write An Algebraic expression to show the total amount he expects to get paid if he gets his bonus after M months
Answer by stanbon(75887)

Alex earns $55,000 as a salesman. he also earns $100 per a month if he reaches his sales goal. Write An Algebraic expression to show the total amount he expects to get paid if he gets his bonus after M months
-----
Each month his basic salary earns him 55000/12 = $4583.33
Equation:
A(M) = (55000/12)M + 100M = $4683.33M
=======================================
Cheers,
Stan H.